How am I supposed to work out this equatin? I hve tried numeros times, but I just don't understnd it.

-20-(-50)=x/9

Answer by Alan3354(69443)

-20-(-50)=x/9
-20 + 50 = x/9
30 = x/9
270 = x
x = 270